OpenStudy (anonymous):

what is the name of the shape graphed by the function theta= 5pi/3 (a) Limacon with inner loop (b) spiral (c)circle (d) line

OpenStudy (anonymous):

In all but one of those, to draw it on a piece of paper you have to adjust your angle.

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

You can do this to check. Just plot the function. Plot the points where the angle is constant on the polar plane.

OpenStudy (anonymous):

but 5pi/3 is the only thing given

OpenStudy (anonymous):

 When you have a constant angle like the given 5pi/3 you get a line on the polar plane You can also get an idea by moving back to the cartesian plane Assume theta is given theta = tan(y/x) arctan(theta)=y/x y=arctan(theta)*x ===> a line